OpenStudy (isuckatmath):

I have eoct tomorrow. i don't get the quadratic formula. the question is, "solve the equation using the quadratic formula" x2 + 3x - 7 = 0 What's a,b, and c?

OpenStudy (anonymous):

The quadratic formula is: \[ax^2 + bx + c = 0 \implies x = \frac{-(b) \pm \sqrt{(b)^2 - 4(a)(c)}}{2(a)}\]

OpenStudy (anonymous):

So looking at your equation: \[x^2 + 3x - 7 = 0\] What number corresponds to a, b, and c?

OpenStudy (isuckatmath):

x^2=a 3x=b 7=c?

OpenStudy (anonymous):

Ok, so If I plug in those values for a, b, and c in the left part of my formula I should get the equation you have.

OpenStudy (anonymous):

\[(x^2)x^2 + (3x)x + (7) = 0\] Is that your equation?

OpenStudy (anonymous):

Still thinking?

OpenStudy (isuckatmath):

no theres only 1 x^2 x2 + 3x - 7 = 0 thats my equation

OpenStudy (anonymous):

right. So what do we need to put in the blanks: \[( )x^2 + ()x + () = 0\] To get something that is equal to your equation?

OpenStudy (anonymous):

Err equivalent rather.